ulfur trioxide (SO3) is produced from the oxidation of SO2 as follows:


2 SO2 + O2 → 2 SO3


A 23-g sample of SO2 gives 18 g of SO3. Calculate the percent yield of sulfur trioxide, SO3.


A. 10 %


1
Expert's answer
2022-06-03T05:33:04-0400

m(SO3) = (m(SO2)*2Mr(SO3)) / 2Mr(SO2)

m(SO3) = (23 g * 2*80 g/mol) / 2*64 g/mol = 28.75 g


percent yield  = (18 g / 28.75 g) * 100% = 62.62 %